# Towards Improved Text-Aligned Codebook Learning: Multi-Hierarchical Codebook-Text Alignment with Long Text

### v1

@@ 1 Introduction
However, the augmented text usually contains hundreds of words, making it challenging for pre-trained language models [43, 9] to capture and retain all key semantic information during encoding. Additionally, encoding and using such lengthy text imposes a significant computational burden. To address this issue, we propose splitting the long text into multiple granularities for encoding, i.e., word, phrase, and sentence. Its advantage is that the long text semantics can be fully encoded without losing any key semantic knowledge by resorting to different granularities. Unfortunately, we are in another dilemma. The multi-granularity text semantics and single code sequence semantics are structurally inconsistent, making direct alignment more difficult. To handle this, we propose a hierarchical encoder that encodes images into multi-hierarchical code representations, each hierarchy corresponding to a specific semantic granularity text. This design ensures a more precise and structurally consistent alignment relationship between image codes and multi-granularity text semantics. Based on the consistent alignment relationship, we propose a novel sampling-based alignment strategy, which can achieve fine-grained codebook-text alignment while not introducing much computational overhead. As a result, a novel text-augmented codebook learning framework, termed Text-Augmented VQ (TA-VQ), is proposed for improved text-aligned codebook learning.
@@
  • We propose to encode long text semantics at three granularities, i.e., word, phrase, and sentence, aiming to retain all key semantic knowledge. Resorting to this multi-granularity text semantics, a hierarchical encoder and novel sampling-based alignment strategy are designed to achieve fine-grained codebook-text alignment.
@@ 4 Proposed Method: TA-VQ
Step 2: Multi-Granularity Text Encoding. Although the generated long text can provide richer semantic knowledge, it also brings a key challenge, that is, how to effectively encode the text. To address this, we propose to split the long text into multiple granularities for encoding, i.e. word, phrase, and sentence. Specifically, given a long text $t$, we employ a text process tool (i.e., TextBlob [36]) to extract all sentences and phrases. Following this, the BERT language model [9] is employed to encode them. As a result, the sentence and phrase semantics can be obtained, which is denoted by $t_{s}=\{s_{1},s_{2},\cdots,s_{|t_{s}|}\}$, where $s_{i}\in\mathbb{R}^{d}$ denotes $i$-th sentence representation with dimension $d$, and $t_{p}=\{p_{1},p_{2},\cdots,p_{|t_{p}|}\}$, where $p_{i}\in\mathbb{R}^{d}$ denotes $i$-th phrase representation, respectively. For word granularity, inspired by [60], we focus on extracting visual-relevant words, such as nouns, adjectives, and quantifiers, rather than encoding all words. As a result, the word semantic can be obtained by BERT pre-trained word embedding, which is denoted by $t_{w}=\{w_{1},w_{2},\cdots,w_{|t_{w}|}\}$, where $w_{i}\in\mathbb{R}^{d}$ denotes $i$-th word representation. Resorting to text semantics of different granularities, the long text can be effectively encoded without losing any key semantic knowledge, which is beneficial for achieving fine-grained alignment.
@@ 5.2 Discussion of Results
Table 1 compares our method and the state-of-the-art VQ-based method in image reconstruction performance. From the results, we can obtain several key conclusions: 1) Compared with the backbone network (i.e., VQ-GAN, and CVQ), our method can learn more robust codebook representation by introducing pre-trained text semantics to significantly improve reconstruction performance, indicating the effectiveness of our method; 2) Our method surpasses LG-VQ across all datasets, largely due to its ability to integrate long text with richer semantic knowledge and effectively encode them at different granularities to learn more robust codebook learning; 3) Compared to VQCT, our method can achieve lower reconstruction errors when using the CVQ backbone, which suggests the effectiveness and generality of method. Notably, our method achieves superior reconstruction and significantly enhances codebook performance across various downstream tasks, suggesting our method’s effectiveness. Further qualitative comparisons are provided in the supplementary materials.

### v2

@@ 1 Introduction
However, the augmented text usually contains hundreds of words, making it challenging for pre-trained language models [43, 9] to capture and retain all key semantic information during encoding. Additionally, encoding and using such lengthy text imposes a significant computational burden. To address this issue, we propose splitting the long text into multiple granularities for encoding, i.e., word, phrase, and sentence. Its advantage is that the long text semantics can be fully encoded without losing key semantic knowledge by resorting to different granularities. Unfortunately, we are in another dilemma. The multi-granularity text semantics and single code sequence semantics are structurally inconsistent, making direct alignment more difficult. To handle this, we propose a hierarchical encoder that encodes images into multi-hierarchical code representations, each hierarchy corresponding to a specific semantic granularity text. This design ensures a more precise and structurally consistent alignment relationship between image codes and multi-granularity text semantics. Based on the consistent alignment relationship, we propose a novel sampling-based alignment strategy, which can achieve fine-grained codebook-text alignment while not introducing much computational overhead. As a result, a novel text-augmented codebook learning framework, termed Text-Augmented VQ (TA-VQ), is proposed for improved text-aligned codebook learning.
@@
  • We propose to encode long text semantics at three granularities, i.e., word, phrase, and sentence, aiming to retain as much key semantic knowledge as possible. Resorting to this multi-granularity text semantics, a hierarchical encoder and novel sampling-based alignment strategy are designed to achieve fine-grained codebook-text alignment.
@@ 4 Proposed Method: TA-VQ
Step 2: Multi-Granularity Text Encoding. Although the generated long text can provide richer semantic knowledge, it also brings a key challenge, that is, how to effectively encode the text. To address this, we propose to split the long text into multiple granularities for encoding, i.e. word, phrase, and sentence. Specifically, given a long text $t$, we employ a text process tool (i.e., TextBlob [36]) to extract all sentences and phrases. Following this, the BERT language model [9] is employed to encode them. As a result, the sentence and phrase semantics can be obtained, which is denoted by $t_{s}=\{s_{1},s_{2},\cdots,s_{|t_{s}|}\}$, where $s_{i}\in\mathbb{R}^{d}$ denotes $i$-th sentence representation with dimension $d$, and $t_{p}=\{p_{1},p_{2},\cdots,p_{|t_{p}|}\}$, where $p_{i}\in\mathbb{R}^{d}$ denotes $i$-th phrase representation, respectively. For word granularity, inspired by [60], we focus on extracting visual-relevant words, such as nouns, adjectives, and quantifiers, rather than encoding all words. As a result, the word semantic can be obtained by BERT pre-trained word embedding, which is denoted by $t_{w}=\{w_{1},w_{2},\cdots,w_{|t_{w}|}\}$, where $w_{i}\in\mathbb{R}^{d}$ denotes $i$-th word representation. Resorting to text semantics of different granularities, the long text can be effectively encoded without losing key semantic knowledge, which is beneficial for achieving fine-grained alignment.
@@ 5.2 Discussion of Results
Tab. 1 compares our method and the state-of-the-art VQ-based method in image reconstruction performance. From the results, we can obtain several key conclusions: 1) Compared with the backbone network (i.e., VQ-GAN, and CVQ), our method can learn more robust codebook representation by introducing pre-trained text semantics to significantly improve reconstruction performance, indicating the effectiveness of our method; 2) Our method surpasses LG-VQ across all datasets, largely due to its ability to integrate long text with richer semantic knowledge and effectively encode them at different granularities to learn more robust codebook learning; 3) Compared to VQCT, our method can achieve lower reconstruction errors when using the CVQ backbone, which suggests the effectiveness and generality of method. Notably, our method achieves superior reconstruction and significantly enhances codebook performance across various downstream tasks, suggesting our method’s effectiveness. Further qualitative comparisons are provided in the supplementary materials.
@@ 6 Conclusions
Acknowledgement
This work was supported by the Guangdong Basic and Applied Basic Research Foundation under Grant No. 2025A1515011674, the Shenzhen Peacock Program under Grant No. ZX20230597, NSFC under Grant No. 62272130 and Grant No. 62376072. It was also supported by the Major Key Project of Peng Cheng Laboratory.
Supplementary Material
7 Experiment Details
Image Generation: For semantic image synthesis, unconditional generation, and image completion, we follow the default setting of VQ-GAN-Transformer [^1] https://github.com/CompVis/taming-transformers. We use a 16-layer transformer as the backbone network with the head of 16 and dimension of 1024. More settings can be found in [13]. For the text-to-image task, we follow the default setting of VQ-Diffusion [^2] https://github.com/microsoft/VQ-Diffusion. We use a 19-layer transformer as the backbone network with the head of 16 and dimension of 1024. The diffusion step is 100. More settings can be found in [17].
Visual Grounding: The RefCOCO dataset [59] comprises 19,994 images with 50,000 referred objects, each associated with multiple referring expressions, totaling 142,210 expressions. Following [29], we adopt the RefCOCOgumd [40] protocol to partition the dataset, resulting in 42,404 expressions for training, 3,811 for validation, and 3,785 for testing. We follow the default setting of LG-VQ [29], we first apply the VQ model to quantize the image into discrete token representations, which are then processed by a learnable adapter network (e.g., a 2-layer MLP). We use a pre-trained CLIP model [43] to encode the object descriptions, and then concatenate the image token representations with the text embeddings. These concatenated features are fed into a trainable transformer, and the last hidden embeddings are used to predict the object box. We use a 3-layer transformer as the backbone network with the head of 16 and dimension of 512. The learnable adapter network is 2-layer MLP with ReLU activation function. The prediction network of the object box is 3-layer MLP with ReLU activation function.
Visual Text Reasoning: The COCO-QA dataset [46] is automatically generated from captions in the Microsoft COCO dataset [30] and contains 78,736 training questions and 38,948 test questions, based on 8,000 and 4,000 images, respectively. The dataset comprises four types of questions: object (70%), number (7%), color (17%), and location (6%). Each answer is a single-word. We follow the default setting of LG-VQ [29]. we first use the VQ model to quantize the image into discrete token representations, and then feed it into a learnable Adapter network (e.g., 2-layer MLP). Following this, we concatenate image token representation with the text embedding and feed them into a pre-trained language model (i.e., GPT2 [44]). We adjust the output of the language model to adapt to different tasks. The learnable adapter network is 2-layer MLP with ReLU activation function. For the image captioning task, we generate captions by predicting the next token in an autoregressive manner. For the VQA task, we feed the last hidden embedding into 2-layer MLP to predict the answer.
Figure 8: Visualization of codebook-word-phrase alignment.
Table C: Ablation study of $f_{j}$ on CUB.
[8, 16, 32] [2, 8, 16] [2, 4, 16] [16, 16, 16] [4, 8, 16]
[…]
8 More Ablation Study
Can our codebook be effectively aligned with the text? In Fig. 8, we visualize an example to demonstrate that our method achieves satisfactory alignment.
Varying the hierarchical grid features $f_{j}$. We provide the impact of different $f_{j}$ on performance in Tab. C.
9 Examples of Origin Caption and Long Text
We provide some examples of origin caption and long text in Fig. 9 for MS-COCO dataset, in Fig. 10 for CelebA-HQ dataset, and in Fig. 11 for CUB-200 dataset.
Figure 9: Examples of origin caption and long text on MS-COCO dataset.
Figure 10: Examples of origin caption and long text on CelebA-HQ dataset.
Figure 11: Examples of origin caption and long text on CUB-200 dataset.
10 More Examples and Qualitative Results
We provide qualitative comparison of image reconstruction in Fig. 12, unconditional generation in Fig. 13, image completion in Fig. 14, semantic image synthesis in Fig. 15, text-to-image in Fig. 16, visual grounding in Fig. 17, and VQA in Fig. 18.
Figure 12: Reconstructions from different models. The red-color boxes highlight reconstruction details.
Figure 13: Visualizations for unconditional generation on CelebA-HQ.
Figure 14: Visualizations for image completion on CelebA-HQ.
Figure 15: Visualizations for semantic synthesis on CelebA-HQ.
Figure 16: Visualizations for text-to-image on CelebA-HQ.
Figure 17: Visualizations for visual grounding. Blue boxes are the ground-truth, red boxes are the model predictions.
Figure 18: Visualizations for visual question answering (VQA).